OpenStudy (anonymous):

Find the slope and y intercept for the given line. 3 x + 2 y = 7 m = __________ y -intercept __________ Find the slope and y intercept for the given line. 3 x + 2 y = 7 m = __________ y -intercept __________ @Mathematics

OpenStudy (chriss):

solve for y and your coefficient for x is your slope, and the constant will be the y intercept. \[y=(-3/2)x+7/2\] -3/2=m 7/2=y-intercept

OpenStudy (anonymous):

Oh Ok. Can you help me with another?

OpenStudy (anonymous):

Input the equation of the given line in standard form.The line with m = -3/4 and b = -2

OpenStudy (chriss):

Yeah

OpenStudy (chriss):

y=(-3/4)x-2

OpenStudy (chriss):

I assume the standard form you mean is y=mx+b

OpenStudy (anonymous):

Yeah I think so... No I think it is Ax+By=c

OpenStudy (chriss):

sorry, I was having a connection issue... ok, you want to start with the y=mx+b for I gave you above \[y=(-3/4)x-2\] then multiply 4 by both sides 4y=-3x-8 then add 3x to both sides 3x+4y=-8

OpenStudy (anonymous):

Thank-you so much,

OpenStudy (chriss):

you are welcome :)

OpenStudy (anonymous):

Find the slope of the following line. y - 3 = 0 m = Is this o,3?

OpenStudy (chriss):

this is a line at y=3. your slope is going to be 0.

OpenStudy (anonymous):

Then I was close?

OpenStudy (chriss):

one way to think of it is if you solve it for y y=0+3 0,3 is a coordinate, and your question asked for a slope  Anyway, I'm terrible at drawing on this thing apparently, but the graph is what you are looking at. the top line is at y=3 and as you can see it is horizontal... a horizontal line has a slope of 0.

OpenStudy (anonymous):

So then y=3

OpenStudy (anonymous):

OR m=3

OpenStudy (chriss):

y=3 m=0 your question asked for the slope. If we go back to the info you were given and put it in the form y=mx+b you get y=0x+3

OpenStudy (anonymous):

Ok, Hey I have another one I got the answer but am unsure. Find the slope of the following line. 3 y + 2 x = 1 m = -2/3x

OpenStudy (chriss):

almost... m=-2/3 the x is the x-coordinate, and is not part of the slope

OpenStudy (anonymous):

Thanks... Find the slope of the following line. 2 x - 1 = 0 m = ( I put undefined) Not sure why it's wrong

OpenStudy (chriss):

your slope should be 2. anytime you see a coefficient for x, that should tip you off that that number is your slope. The equation for this line is y=2x-1. Setting y to 0 (as was done here) you solve for x to find that x=1/2 In other words you have a line defined by y=2x-1 and it passes through the point (1/2,0)

OpenStudy (anonymous):

Oh that makes sense.

OpenStudy (anonymous):

It says that I am wrong.

OpenStudy (chriss):

It says that 2 is not the slope?

OpenStudy (anonymous):

Yeah it does... I do school work online. And it is really hard to do it without a teacher.

OpenStudy (chriss):

I hear ya. I'm not sure why it is telling you that. The answer I gave you is the correct slope. Just flat out, once you get a problem down to y=mx+b form, the m is going to be your slope.

OpenStudy (anonymous):

Well, Ok here is another one. Find the slope of the following line. Leave answers in fractional form. 3 x - 2 y = 4

OpenStudy (anonymous):

I really don't where to start.

OpenStudy (chriss):

let's let you try it first... the hint is that you want to solve it for y.

OpenStudy (anonymous):

Ok So 2y = -3x + 4/2

OpenStudy (chriss):

so what would be the first thing you might want to get rid of to get y by itself? It's going to take at least 2 steps to do it.

OpenStudy (anonymous):